# Energy based diffusion generator for efficient sampling of Boltzmann distributions

### v2

@@ Appendix F Ablation study
For VAE w/ GHD, the decoder is consistent with the structure in Algorithm 1, and the encoder is the same as decribed above. For EDG w/o GHD, we omit the leap-frog component (refer to Line 2-8 in the algorithm) and the decoder is composed of the network $a(\zeta;\phi)$ and a final gaussian part $\mathcal{N}(\mu,\Sigma)$ (refer to Line 1, 9 in the algorithm).
@@
Table 5: The Maximum Mean Discrepancy (MMD) between 5,000 samples generated by each generator and the reference samples.
Mog2 Mog2(i) Mog6 Mog9 Ring Ring5
VAE w/o GHD $1.86$ $1.62$ $2.57$ $2.10$ $0.12$ $0.23$
VAE w/ GHD $\mathbf{0.01}$ $2.43$ $0.15$ $0.59$ $1.68$ $0.63$
EDG w/o GHD $0.06$ $1.01$ $0.04$ $0.05$ $0.02$ $0.04$

### v3

@@ Appendix I Computation cost
For VAE w/ GHD, the decoder is consistent with the structure in Algorithm 3, and the encoder is the same as described above. For EDG w/o GHD, we omit the leap-frog component (refer to Line 2-8 in the algorithm) and the decoder is composed of the network $a(\zeta;\phi)$ and a final gaussian part $\mathcal{N}(\mu,\Sigma)$ (refer to Line 1, 9 in the algorithm).
Table 8: The MMD between 5,000 samples generated by each generator and the reference samples.
Mog2 Mog2(i) Mog6 Mog9 Ring Ring5
VAE w/o GHD $1.86$ $1.62$ $2.57$ $2.10$ $0.12$ $0.23$
VAE w/ GHD $\mathbf{0.01}$ $2.43$ $0.15$ $0.59$ $1.68$ $0.63$
EDG w/o GHD $0.06$ $1.01$ $0.04$ $0.05$ $0.02$ $0.04$
EDG $\mathbf{0.01}$ $\mathbf{0.50}$ $\mathbf{0.01}$ $\mathbf{0.02}$ $\mathbf{0.01}$ $\mathbf{0.02}$
@@
The step size of Decoder: Step sizes will impact the model’s performance for different tasks, selecting appropriate step sizes is crucial to achieve optimal results. To mitigate the influence of step size on the model’s effectiveness, we treat them as trainable parameters with a hyperparameter $\epsilon_{0}$ in Sec. 3.4.2). In our experiments, we systematically evaluated $\epsilon_{0}$ from 0.005 to 0.1 in increments of 0.005, selecting the optimal value based on training performance. The optimal value is provided in Tab. 9.
Table 9: Step size hyperparameters across tasks
Task 2d energy except Rings 2d-Rings Bayesian Regression
$\epsilon_{0}$ 0.1 0.03 0.05
Task Bayesian Regression-Cover LJ13 LJ15 Ising model
$\epsilon_{0}$ 0.005 0.03 0.015 0.1
Network architecture: To evaluate the effectiveness of the hierarchical design, we compare it with an EDG model using standard fully connected linear layers. As shown in Fig. 10, the hierarchical architecture better captures the underlying structure of the data and achieves superior performance. In future work, we plan to design specific network architectures for different tasks.
Figure 10: The $\log Z$ estimation curves under different model architectures in the Ising model task at $T=2.0$. Hierarchy denotes the model architecture in H and MLP refers to EDG models where all networks are implemented as 3-layer MLPs with 32 hidden units and ReLU activation functions.